Please help me solve this problem: ((( 3 more than twice a number is 2 ))) I have to write this as an algebraic expression or equation.
Answer by scoop(27)

The number will be x
So three more than twice a number is
2x+3=2
then subtract the three from both sides to get like terms
2x=2-3
2x=-1
divide both sides by the 2 to get x by itself
x=-1/2 (your answer)
to check this answer multiply
2(-1/2) +3 =2
-2/2 +3=2
-1+3=2